OpenStudy (anonymous):

Sam's plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

\((14+x)\cdot(18+x)=480\), solve for \(x\).

OpenStudy (anonymous):

the are of the tv is \[14\times 18=252\] let x be the trim 252+x=480 x=480-252 x=225

OpenStudy (anonymous):

@babsarica That gives the area of the trim, but the question was asking for the width of the trim.

OpenStudy (anonymous):

sorry x=228 then 252-228=24 therefore the trim is 24 in wide

OpenStudy (anonymous):

Not quite sure what you're doing there, but that's incorrect. The equation that I originally replied with is the correct method to solve the problem. Solving for \(x\) gives you 6, which is the correct width of the trim.